# Isbell’s subfactor projections in a noetherian form

### v1

@@ 7 Isbell’s results on subfactors
by the Butterfly Lemma (Lemma 7.4), $[Y_{j+1}^{i},Y_{j}^{i}]$ and $[X_{i+1}^{j},X_{i}^{j}]$ project onto each other and hence, in particular, $X_{i+1}^{j}\neq X_{i}^{j}$ and so $[X_{i+1}^{j},X_{i}^{j}]$ is a subfactor in the second subnormal series. So the two subnormal series are projectively isomorphic.
8 Counterexample to Isbell’s claim on coarsest refinements
@@
    Every subfactor $[X^{-},X^{+}]$ contained in some $[X_{i+1},X_{i}]$ and onto which some subfactor smaller than $[Y_{j+1},Y_{j}]$ projects is contained in the projection $[Y_{j+1},Y_{j}][X_{i+1},X_{i}]$.
@@
unlike what is claimed in (E1). Now, to contradict (E2), we first note that the projectively isomorphic refinements produced by the two given subnormal series, using the contruction described in Theorem 7.4, are both the series $Y_{0}\supset Y_{1}\supset Y_{2}$. However, this pair of projectively isomorphic refinements is not the coarsest pair. The subnormal series
@@
{references*}

### v2

@@ 3. Isbell’s results on subfactors
by the Butterfly Lemma (Lemma 3.3), $[Y_{j+1}^{i},Y_{j}^{i}]$ and $[X_{i+1}^{j},X_{i}^{j}]$ project onto each other and hence, in particular, $X_{i+1}^{j}\neq X_{i}^{j}$ and so $[X_{i+1}^{j},X_{i}^{j}]$ is a subfactor in the second subnormal series. So the two subnormal series are projectively isomorphic. ∎
4. Counterexample to Isbell’s claim on coarsest refinements
@@
    Every subfactor $[X^{-},X^{+}]$ contained in some $[X_{i+1},X_{i}]$ and onto which some subfactor smaller than $[Y_{j+1},Y_{j}]$ projects is contained in the projection
    $[Y_{j+1},Y_{j}][X_{i+1},X_{i}].$
@@
unlike what is claimed in (E1). Now, to contradict (E2), we first note that the projectively isomorphic refinements produced by the two given subnormal series, using the construction described in Theorem 3.5, are both the series $Y_{0}\supset Y_{1}\supset Y_{2}$. However, this pair of projectively isomorphic refinements is not the coarsest pair. The subnormal series
@@
The proof of (E1) and (E2) in [12] is a single sentence claiming that the following implication holds:
$[Y^{-},Y^{+}]\subseteq[Y_{j+1},Y_{j}]\Rightarrow[Y^{-},Y^{+}][X_{i+1},X_{i}]\subseteq[Y_{j+1},Y_{j}][X_{i+1},X_{i}].$
While this claim holds indeed, it does not imply (E1), since $[Y^{-},Y^{+}][X_{i+1},X_{i}]$ need not be the same as $[X^{-},X^{+}]$ in (E1) (where $[Y^{-},Y^{+}]$ is the subfactor contained in $[Y_{j+1},Y_{j}]$ which projects onto $[X^{-},X^{+}]$). Indeed, in the counterexample above,
$[Y^{-},Y^{+}][X_{i+1},X_{i}]=[Y_{2},Y_{1}][X_{1},X_{0}]=[Y_{2},Y_{1}]\neq[X^{-},X^{+}].$
5. On other abstract approaches to the butterfly lemma
The butterfly lemma and the refinement theorem have been abstracted to numerous general contexts. For instance, already the book on group theory by Zassenhaus [28] (originally published in 1958), discusses an abstraction in general lattices. Our context is more nuanced in that, firstly, we care about constructing the actual isomorphisms between groups instead of just establishing correspondences inside a subgroup lattice and secondly, because of the self-dual context, not all subgroups are necessarily conormal in our context. The first of these nuances requires an interplay between a lattice-theoretic context and a categorical context. The context of a $\gamma$-category of Burgin [4], for instance, provides such interplay. As remarked in [2], this context brings together semi-abelian categories and Puppe-Mitchell exact categories [21, 19] (which are the same as pointed Grandis exact categories). It is evident from the properties of these categories discussed in [2] that subobjects in a $\gamma$-category constitute a noetherian form of the category if and only if they admit binary joins (a requirement fulfilled in the case of semi-abelian categories, but which is also expected to be fulfilled in the standard non-semi-abelian examples, such as the category of torsion-free abelian groups mentioned in [2], for instance). In this noetherian form, all subgroups will be conormal. So up to the mild limitation of existence of joins of subobjects, the Jordan-Hölder Theorem obtained in [2] could be deduced from our refinement theorem as a consequence. It should be noted, however, that the approach in the proof of the Jordan-Hölder Theorem in [2] is different to a classical approach that passes via the refinement theorem.
Two different but similar approaches in abstracting the butterfly lemma are given in [27] and [22], which have inspired an abstraction in [20] to a fairly general framework based on the notion of a star-regular category [9]. The contexts subsumed here include all regular categories [1] having certain pushouts (the case considered in [22]). In this case, the role of subgroups and the subgroups normal to those are played by entities of different types, which makes it unclear how subnormal series could be defined and hence whether a refinement theorem can be established. It is of course nevertheless interesting that a form of the butterfly lemma can be established in cocomplete regular categories. The class of cocomplete regular categories is significantly wider than the class of semi-abelian categories. In fact, it includes all quasi-varieties of universal algebras. As shown in [15], the formula $f^{-1}f(X)=X\vee\mathsf{Ker}f$ of Axiom 2.2 holds for the form of subobjects in a regular category if and only if it is protomodular [3]. So the form of subobjects in a (cocomplete) regular category need not be noetherian. This does not exclude the possibility of there being another form over a (cocomplete) regular category that is noetherian.
The context of [27] is better comparable to that of a noetherian form. The refinement theorem has been obtained in this context under an additional mild axiom in [7]. The axioms in [27, 7] are fairly weak and hold, for instance, in ideal-determined categories [14], which is a wider class of categories than the class of semi-abelian categories: an ideal determined category is semi-abelian if and only if it is protomodular. The form of subobjects of an ideal-determined category misses to be noetherian precisely because of the failure of the formula $f^{-1}f(X)=X\vee\mathsf{Ker}f$ of Axiom 2.2. The reason why this formula is not needed for the approach to the butterfly lemma in [27] is that there, already in the formulation of the butterfly lemma, Wyler replaces the join $S\vee N$ of subgroups, when $N$ is a normal subgroup, with the construction $\pi_{N}^{-1}\pi_{N}(S)$.
While in some ways the context of a noetherian form is more restrictive than the categorical contexts discussed above, in other ways it is less restrictive: the contexts of [27, 4, 2], for instance, require pointedness of the category, while the contexts of [22, 20] require existence of finite limits (and moreover, regularity of the category). We should also point out that the abstract approaches to the butterfly lemma and the refinement theorem discussed above do not establish these results in their full strength as formulated by Isbell [12], which was the main goal of the present paper.